4. Evaluate Sl y(+12) ds, where S = {(1,y,z) : 2 = 4 – r?,05152,0 <ys 5).
Answer #1

02704 5 sy ( 2² + XZ) ds S= {(113,2): 2-4-43, 0224 24, oty = 5} s j a {2+ x (4-12)} dady 20 5 24 2X-7=z نح s x=0 y=o 24 5 7:2## Dont FORGET to give RATINGS.